# Automating Legal Concept Interpretation with LLMs: Retrieval, Generation, and Evaluation

### v2

@@ 5.6 Result
Table 1: Main results of automatic evaluation on the Legal Concept Entailment task, the best is bolded and the second is underlinded. CS represents the consistency score. We use Qwen2.5-72B to generate concept interpretations and employ Qwen2.5-72B/14B to perform the LCE task.
We report the performance of our method and all baselines on the LCE Task in Table 1. Overall, ATRIE achieves the best performance across nearly all models and evaluation metrics, showcasing the effectiveness of our framework and the necessity of its core components.
@@ 5.6.1 Classification Task
(1) LLMs possess some level of discriminative ability.
The performance of "w/o Interpretation" surpasses that of random guessing. Besides, CoT’s performance surpasses that of Zero-shot, demonstrating that step-by-step reasoning benefits the LCE Task.
@@
"w/ Interpretation" significantly outperforms "w/o Interpretation." "w/ Direct Interpretation" shows that LLMs can leverage their extensive internal knowledge to reason about vague concepts and generate useful legal concept interpretations. "w/ Judicial Interpretation" falls short of "w/ Direct Interpretation." We attribute this to the relatively simple explanations provided in judicial interpretations, which lack the depth required to guide LLMs in evaluating the applicability of vague concepts to specific cases. The performance of "w/ Expert Interpretation" is inferior to ATRIE. We attribute this to the fact that expert-written interpretations are often overly abstract and detailed, which results in poorer readability. We will further discuss this in the human evaluation (Sec 7).
(3) Utilizing relevant cases is necessary.
@@ 6.1 How to retrieve cases
To verify the importance of each step in our process of retrieving relevant cases, we compare the performance of the following settings: (1) No retrieval, where no cases are retrieved (i.e., LLM Direct Interpretation); (2) String Match, which does not use LLM to filter cases or perform label balancing; (3) String Match + Filter, which does not perform label balancing. We ensure that the number of cases retrieved by each method is consistent. Table 2 shows that every component of our retrieval method is necessary.
@@ 6.2 Number of cases
We investigated the impact of using different numbers of case judgments on the quality of generated concept interpretations. Specifically, we sampled different numbers of reasons from the extracted reason set $\mathcal{R}$ as input to the LLM. Figure 4 shows that more input reasons lead to higher-quality interpretations.
The more cases legal experts review, the more comprehensive their concept interpretations become. Our findings align with legal experts’ experiences, showcasing LLMs’ ability to analyze numerous cases effectively and highlighting their advantage in aiding legal concept interpretation.
@@ 6.3 Which parts of a case are useful?
In Section 4.2, we only extract a few sentences discussing the concept from the court view of each relevant case without including the complete fact description and court view. We aim to investigate whether this might result in the loss of important information from the case, potentially affecting the generation of interpretations. To explore this, we compared three different approaches to representing the relevant information of a case during the interpretation generation step: (1) Court View: the part of the judgment where the judge explains the legal rationale and interprets the basis of the ruling; (2) Summarized Fact and Court View: The facts section in case judgments is often lengthy and contains excessive detail. To address this, we first use an LLM to summarize the facts and then concatenate it with the court view section; (3) Extracted Reason: Extracted reasons in Section 4.2.
@@
Table 3: Results of using different parts of case judgment to generate interpretations.
@@ 6.5 Are legal LLMs more effective?
We also use legal LLMs to generate concept interpretations. ATRIE requires analyzing hundreds of cases, with an average input length of 17k tokens. In contrast, among the currently available Chinese legal LLMs, Farui-plus [^7] https://tongyi.aliyun.com/farui—which offers the longest maximum context length—supports only up to 12k tokens (Appendix C.6). Thus, we restrict the input length to within 10k tokens and compare the concept interpretations generated by Farui-plus and Qwen2.5-72B under identical input conditions. Table 5 shows that general-purpose LLM Qwen significantly outperforms legal LLM Farui in interpreting legal concepts.
@@ 7.1 Evaluation Metrics
We recruited 2 legal experts who have passed China’s Unified Legal Profession Examination to assess the legal concept interpretations generated by Qwen2.5 (72B). They collaboratively establish five evaluation criteria and score the interpretations: (1) Accuracy (Acc.), (2) Informativeness (Info.), (3) Normativity (Norm.), (4) Comprehensiveness (Comp.), (5) Readability (Read.). We use a 10-point Likert scale, where 1 represents "very poor" and 10 represents "very good".  [^8] Details about the metrics and human evaluation are discussed in Appendix E.
@@ 7.2 Results
We compare three different interpretations in Sec 5.5 for each of the 16 legal concepts. In Table 6, we have several observations: (1) The average score of ATRIE is the highest, indicating that our interpreter can generate legal concept interpretations comparable to those produced by legal experts. (2) The Comprehensiveness score of ATRIE is much higher than Expert Interpretation, indicating that having LLMs read a vast number of cases helps generate more comprehensive concept interpretations. (3) Expert Interpretation (EI) receives the lowest score in Readability, indicating that the interpretations written by legal experts tend to be abstract or complex, which hinders understanding by both humans and LLMs. (4) In Accuracy, Informativeness, and Normativity, ATRIE shows improvements over Direct Interpretation (DI). Although there are still minor gaps between ATRIE and Expert Interpretation, it’s important to note that Expert Interpretation was produced by legal experts who spent considerable time.
@@ 8 Conclusion
In this work, we explore the use of LLMs to address a challenging task in the legal field: Legal Concept Interpretation. By emulating the human approach to doctrinal legal research, we propose a fully automatic framework for retrieving concept-related information, interpreting legal concepts, and evaluating the generated interpretations. Both automatic and human evaluations demonstrate that our generated interpretations are useful and comparable to those written by legal experts. Our study suggests considerable potential for using LLMs to assist legal experts in legal interpretation and beyond.
@@
Legal datasets frequently contain sensitive details about individuals and organizations, and improper handling can result in significant privacy violations. To safeguard this information, the case judgment dataset used in our experiments is thoroughly anonymized.
@@
Appendix A The structure of case judgments
A Case Judgment in China can generally be divided into five sections: header, facts, court view, verdict, and conclusion. The header includes the name of the court, the type of document, case number, basic information about the parties involved, the origin of the case, and details about the judicial panel and trial method. The facts section outlines the plaintiff’s claims, facts, arguments, and the defendant’s admissions regarding the plaintiff’s factual assertions. The court view section provides the rationale for the judgment and the legal basis upon which it is made. The verdict contains the decision on substantive issues of the case. Finally, the conclusion ends the judgment document formally.
@@ Appendix C Details of ATRIE
C.1 Implementation details
We filtered 2,642 cases and extracted the same number of reasons for generating concept interpretations. On average, each concept was associated with 165 cases. We use the open-source LLM Qwen2.5-72B-Instruct with a maximum context length of 128k tokens to generate vague concept interpretations. The temperature is set to 0.9 to encourage more diverse outputs. Detailed prompt information can be found in Appendix G.1.4.
To investigate the effectiveness of our generated interpretations in assisting models with different capabilities, we employ Qwen2.5-72B-Instruct and Qwen2.5-14B-Instruct to perform the LCE task.
To reduce the randomness of the output, the temperature of all LLMs for prediction is set to 0, and the generation process is repeated three times. Among the predictions, we select the label $\hat{l}$ that appears most frequently. From the responses associated with $\hat{l}$, one is randomly chosen, and its reason $\hat{r}$ is extracted for consistency scoring. We use gpt-4o-2024-08-06Achiam et al. (2023) to give the consistency score, setting the temperature to 0.
C.2 Manual inspection of the LLM-annotated data
To evaluate the relevance between the LLM-filtered case judgments and the vague concepts, we randomly sampled 20 cases for each concept from $\mathcal{D}$ and manually assessed their relevance to the vague concepts. The results show that over 96% of the cases are indeed relevant to the vague concepts. In addition, manual inspection of 200 extraction results indicates that the accuracy of Qwen2.5-72B-Instruct in labeling the gold label $l$ and the reasoning $r$ are 98% and 94%, respectively.
C.3 Example of gold labels and reasons
@@
C.4 Detailed results
C.4.1 Different models
As shown in Table 9, to validate the generalizability of our method, we utilized different LLMs to generate interpretations and perform automatic evaluations. Due to the cost constraints of APIs, we conducted experiments on a subset of our LCE dataset. Our findings are as follows: (1) Stronger models demonstrate more remarkable ability to generate concept interpretations. The interpretations generated using Qwen2.5 (72B) and GPT-4o lead to noticeably higher performance improvements than using GPT-4o-mini. (2) Generated concept interpretations can assist even weaker LLMs in accurately understanding vague concepts. In our method, the performance gap between GLM and the other models is significantly smaller than that observed in the Zero-Shot baseline.
@@ C.4.2 Model bias
C.5 Open-source LLMs are also good evaluators
@@
C.6 Why don’t we use legal LLMs in our interpreter?

### v3

@@ 5.5 Baselines
5.6 Implementation details
After filtering, we obtained 2,642 cases and extracted the same number of reasons for generating concept interpretations. On average, each concept was associated with 165 cases. We use the open-source LLM Qwen2.5-72B-Instruct with a maximum context length of 128k tokens to generate vague concept interpretations. The temperature is set to 0.9 to encourage more diverse outputs. Detailed prompt information can be found in Appendix G.1.4.
To investigate the effectiveness of our generated interpretations in assisting models with different capabilities, we employ Qwen2.5-72B-Instruct and Qwen2.5-14B-Instruct to perform the LCE task. We use gpt-4o-2024-08-06Achiam et al. (2023) to give the consistency score. To reduce the randomness of the output, the temperature of all LLMs in LCE task is set to 0.
@@ 5.7 Result
Table 1: Main results of automated evaluation on the Legal Concept Entailment task, the best is bolded and the second is underlinded. CS represents the consistency score. We use Qwen2.5-72B to generate concept interpretations and employ Qwen2.5-72B/14B to perform the LCE task.
We report the performance of our method and all baselines on the LCE Task in Table 1. Overall, ATRIE achieves the best performance across nearly all models, showcasing the effectiveness of our framework and the necessity of its core components.
@@ 5.7.1 Classification Task
(1) LLMs exhibit a certain degree of discriminative ability by leveraging their internal knowledge.
The performance of the "w/o Interpretation" setting surpasses that of random guessing, yet there remains significant room for improvement. This suggests that using the LLM’s performance on the LCE task as a proxy for interpretation quality is a feasible approach.
@@
"w/ Direct Interpretation" outperforms all the "w/o Interpretation" settings, showing that LLMs can leverage their extensive internal knowledge to reason about vague concepts and generate useful legal concept interpretations. "w/ Judicial Interpretation" falls short of "w/ Direct Interpretation." We attribute this to the relatively simple explanations provided in judicial interpretations, which lack the depth required to guide LLMs in evaluating the applicability of vague concepts to specific cases. The performance of "w/ Expert Interpretation" is inferior to ATRIE. We attribute this to the fact that expert-written interpretations are often overly abstract and detailed, which results in poorer readability. We will further discuss this in the human evaluation (Sec 7).
(3) Utilizing relevant cases is necessary for good concept interpretations.
@@ 6.1 Retrieval Method
To verify the importance of each step in our process of retrieving relevant cases, we compare the performance of the following settings: (1) No retrieval does not retrieve cases (i.e., LLM Direct Interpretation); (2) String Match does not use LLM to filter cases or perform label balancing; (3) String Match + Filter only removes the label balancing stage. We ensure that the number of cases retrieved by each method is consistent. Table 2 shows that every component of our retrieval method is necessary.
@@ 6.2 Number of Cases
We investigated the impact of using different numbers of cases on the quality of generated concept interpretations. Specifically, we sampled different numbers of reasons from the extracted reason set $\mathcal{R}$ as input to the LLM. Figure 4 shows that more input reasons lead to higher-quality interpretations.
The more cases legal experts review, the more comprehensive their concept interpretations become. Our findings align with legal experts’ experiences, showcasing LLMs’ ability to analyze numerous cases effectively and highlighting their advantage in aiding legal interpretation.
@@ 6.3 Which parts of a case are useful?
In Section 4.2, we only extract a few sentences discussing the concept from the court view of each relevant case without including the complete fact description and court view. We aim to investigate whether this might result in the loss of important information from the case, potentially affecting the generation of interpretations. To explore this, we compared three different approaches to representing the relevant information in a case during the interpretation generation step: (1) Court View: the part of the judgment where the judge explains the legal rationale and interprets the basis of the ruling; (2) Summarized Fact and Court View: The fact section in a case is often lengthy and contains excessive detail. To address this, we first use Qwen2.5-72B-Instruct to summarize the facts and then concatenate it with the court view section; (3) Extracted Reason: Extracted reasons in Section 4.2.
@@
Table 3: Results of using different parts of cases to generate interpretations.
@@ 6.5 Are Legal LLMs More Effective?
Previously, we utilized general-purpose LLMs for generating legal interpretations. However, we also aimed to evaluate the performance of legal LLMs. ATRIE requires analyzing hundreds of cases, with an average input length of 17k tokens. In contrast, among the currently available Chinese legal LLMs, Farui-plus [^7] https://tongyi.aliyun.com/farui—which offers the longest maximum context length—supports only up to 12k tokens (Appendix C.5). Thus, we restrict the input length to within 10k tokens and compare the concept interpretations generated by Farui-plus and Qwen2.5-72B-Instruct under identical input conditions. Table 5 shows that general-purpose LLM Qwen significantly outperforms legal LLM Farui in interpreting legal concepts. This phenomenon may be attributable to the superior long-text comprehension and generation capabilities of Qwen2.5-72B-Instruct, which enables more effective summarization from extensive texts.
@@ 7.1 Evaluation Metrics
We recruited 2 legal experts who have passed China’s Unified Legal Profession Examination to assess the legal concept interpretations generated by Qwen2.5-72B-Instruct. They collaboratively establish five evaluation criteria and score the interpretations: (1) Accuracy (Acc.), (2) Informativeness (Info.), (3) Normativity (Norm.), (4) Comprehensiveness (Comp.), (5) Readability (Read.). We use a 10-point Likert scale, where 1 represents "very poor" and 10 represents "very good".  [^8] Details about the metrics and human evaluation are discussed in Appendix E.
@@ 7.2 Results
We compare three different interpretations in Sec 5.5 for each of the 16 legal concepts. In the Appendix E.3, we present a case study to illustrate a comparative analysis of the scores for interpretations generated by two distinct methods, along with the rationale behind these scores. In Table 6, we have several observations: (1) The average score of ATRIE is the highest, indicating that our interpreter can generate legal concept interpretations comparable to those produced by legal experts. (2) The Comprehensiveness score of ATRIE is much higher than Expert Interpretation, indicating that having LLMs read a vast number of cases helps generate more comprehensive concept interpretations. (3) Expert Interpretation (EI) receives the lowest score in Readability, indicating that the interpretations written by legal experts tend to be abstract or complex, which may hinder understanding by both humans and LLMs. (4) In Accuracy, Informativeness, and Normativity, ATRIE shows improvements over Direct Interpretation (DI). Although there are still minor gaps between ATRIE and Expert Interpretation, it’s important to note that Expert Interpretation was produced by legal experts who spent considerable time.
@@ 8 Conclusion
In this work, we explore the use of LLMs to address a challenging task in the legal field: Legal Interpretation. By emulating doctrinal legal research, we propose a fully automated framework for retrieving concept-related information, interpreting legal concepts, and evaluating the generated interpretations. Both automated and human evaluations demonstrate that our generated interpretations are useful and comparable to those written by legal experts. Our study suggests considerable potential for using LLMs to assist legal experts in legal interpretation and beyond.
@@
Legal datasets frequently contain sensitive details about individuals and organizations, and improper handling can result in significant privacy violations. To safeguard this information, the case dataset used in our experiments is thoroughly anonymized.
@@
Acknowledgements
This work is supported in part by NSFC (62161160339) and Beijing Science and Technology Program (Z231100007423011). We thank the anonymous reviewers for their valuable suggestions. For any correspondence, please contact Yansong Feng.
Appendix A The structure of cases
A case in China can generally be divided into five sections: header, facts, court view, verdict, and conclusion. The header includes the name of the court, the type of document, case number, basic information about the parties involved, the origin of the case, and details about the judicial panel and trial method. The facts section outlines the plaintiff’s claims, facts, arguments, and the defendant’s admissions regarding the plaintiff’s factual assertions. The court view section provides the rationale for the judgment and the legal basis upon which it is made. The verdict contains the decision on substantive issues of the case. Finally, the conclusion ends the judgment document formally.
@@ Appendix C Details of ATRIE
C.1 Manual inspection of the LLM-annotated data
To evaluate the relevance between the LLM-filtered cases and the vague concepts, we randomly sampled 20 cases for each concept from $\mathcal{D}$ and manually assessed their relevance to the vague concepts. The results show that over 96% of the cases are indeed relevant to the vague concepts. In addition, manual inspection of 200 extraction results indicates that the accuracy of Qwen2.5-72B-Instruct in labeling the gold label $l$ and the reasoning $r$ are 98% and 94%, respectively.
C.2 Example of gold labels and reasons
@@
C.3 Detailed results
C.3.1 Different models
As shown in Table 9, to validate the generalizability of our method, we utilized different LLMs to generate interpretations and perform automated evaluations. Due to the cost constraints of APIs, we conducted experiments on a subset of our LCE dataset. Our findings are as follows: (1) Stronger models demonstrate more remarkable ability to generate concept interpretations. The interpretations generated using Qwen2.5 (72B) and GPT-4o lead to noticeably higher performance improvements than using GPT-4o-mini. (2) Generated concept interpretations can assist even weaker LLMs in accurately understanding vague concepts. In our method, the performance gap between GLM and the other models is significantly smaller than that observed in the Zero-Shot baseline.
@@ C.3.2 Model bias
C.4 Open-source LLMs are also good evaluators
@@
C.5 Why don’t we use legal LLMs in our interpreter?
@@ E.2 Instructions given to annotators
E.3 Case Study
We asked legal experts to conduct a detailed analysis comparing the Direct Interpretation (DI) and our framework ATRIE, which can demonstrate how high-scoring interpretations are concretely superior to low-scoring ones. They both interpret the concept of “serious circumstances” under Article 359 of the Chinese Criminal Law. We believe the difference in scores is a reliable measure of the quality of the interpretations.
  • For Accuracy (8.0 vs. 6.5), ATRIE gives a more legally faithful and analytically precise explanation by explicitly integrating quantifiable benchmarks such as the number of individuals involved and the amount of illegal profit. In contrast, DI does not give detailed criteria, reducing its alignment with actual judicial practice.
  • In Informativeness (7.5 vs. 6.5), ATRIE summarizes a more comprehensive account of various scenarios, incorporating factors such as the means of committing the offense and the duration of the behavior. While they are partially addressed or entirely absent in DI.
  • For Normativity (8.0 vs. 7.0), ATRIE maintains consistent use of legal language and adheres more strictly to legal theory, while DI exhibits a more colloquial tone and lacks binding legal texts, weakening its formal authority.
  • For Comprehensiveness (8.0 vs. 6.5), ATRIE provides a fuller framework by offering both positive and negative case illustrations derived from actual judicial decisions, each structured with clear factual context and legal reasoning. This contrasts with DI’s more abstract and potentially fictional examples.
  • For the hallucination problem, We would like to thank the reviewer for pointing it out, which is really critical in the legal domain. In the case study, we manually verified all the cited cases in ATRIE. All cases are from actual judicial decisions and could be referred to specific sentences in judgment documents. There are no instances of hallucinated or fabricated content. In contrast, the examples in DI appear to be hypothetical or generalized, too abstract to be identified in cases, which adds the risk of hallucination.
E.3.1 Direct Interpretation
1. ****
2. ****
3. ****
  1. ****
  2. ****
  3. ****
  4. ****
  5. ****
  1. ****
  2. ****
  3. ****
  4. ****
  5. ****
E.3.2 ATRIE
1. ****
2. ****
3. ****
  1. ****
     - ****1001000
  2. ****
     - ****10
  3. ****
     - ****125
  4. ****
     - ****247748
  5. ****
     - ****10
     - ****10
  1. ****
     - ****20
  2. ****
  3. ****
  4. ****
  5. ****